OpenStudy (anonymous):

Give the dimensions of a cone and a cylinder that have the same volume. Show that the two solids have the same volume.

OpenStudy (campbell_st):

use the measurements given to ypu and substitute Volume of the cylinder is \[V = \pi r^2 h\] Volume of the cone is \[V = \frac{1}{3} \pi r^2 h\] canculate each volume and compare them. Are they the same ? hope it helps. p.s.. its a poor question given you are told they have the same volume...